# The Atlas Model and SDEs with Boundary Interaction

### v1

@@ 1 Introduction
The Atlas model (1.1) was introduced by Fernholz in his seminal monograph [15] on Stochastic Portfolio Theory (see Example 5.3.3). In this context, the particle states represent the logarithmic market capitalisations of constituents of a large stock market. Since then, there has been a steady interest in finite and infinite versions of the Atlas model [4, 27, 9, 28, 13, 7, 2] and more general systems with coefficients depending on particle ranks. Our goal is to understand the behaviour of the system (1.1) as the number $N$ of particles is taken to infinity. In particular, we wish to derive an SDE that describes the dynamics of a representative particle in the limit $N\to\infty$. Due to the interactive nature of the system, we should expect this SDE to be of McKean–Vlasov type.
@@
for $t\geq 0$. These defining features of the Skorokhod problem in one dimension are recalled in Appendix A, Definition A.1.
@@ 1.1 The Mean-Field Limit
where $\tilde{W}=(\tilde{W}_{t})_{t\geq 0}$ is a Brownian motion and $L^{X^{(2)}-X^{(1)}}=(L^{X^{(2)}-X^{(1)}}_{t})_{t\geq 0}$ denotes the local time at zero of the first gap $X^{(2)}-X^{(1)}$. This formula follows e.g. from [15, Proposition 4.1.11], which derives the dynamics of ranked processes. A priori, it is not at all clear that the exploding drift $\gamma Nt$ on the right-hand side of (1.4) should be offset by the local time at zero of the gap $X^{(2)}-X^{(1)}$ in precisely such a way that we wind up with a continuous and deterministic barrier $b$. This motivates us to search for a relaxation of the reflected McKean–Vlasov SDE (1.3), which make sense even when the limit of $(B^{N})_{N\geq 1}$ could in principle be a much rougher object.
@@
Then, as soon as we can get sufficient control on the magnitude of $B^{N}$ as $N\to\infty$, it follows that $(\beta^{N})_{N\geq 1}$ is tight on $\mathcal{M}_{1}([0,\infty)\times\mathbb{R})$. Here $\mathcal{M}_{1}([0,\infty)\times\mathbb{R})$ is the set of measures $m$ on $0,\infty)\times\mathbb{R}$ such that $m([0,t]\times\mathbb{R})=t$ for $t\geq 0$, topologised with the coarsest topology that makes the maps $\mathcal{M}_{1}([0,\infty)\times\mathbb{R})\to\mathcal{P}([0,n]\times\mathbb{R})$, $m\mapsto\frac{1}{n}m|_{[0,n]\times\mathbb{R}}$ for $n\in\mathbb{N}$, continuous. If the sequence $(\beta^{N})_{N\geq 1}$ is tight on $\mathcal{M}_{1}([0,\infty)\times\mathbb{R})$, it admits a subsequential weak limit $\beta$ with values in $\mathcal{M}_{1}([0,\infty)\times\mathbb{R})$. We wish to phrase our generalisation of McKean–Vlasov SDE (1.3) in terms of such $\beta$. As before, the state $(X_{t})_{t\geq 0}$ of the representative particle takes the form $X_{t}=\xi+W_{t}+L_{t}$ for a nondecreasing continuous process $L=(L_{t})_{t\geq 0}$ started from zero. However, we can no longer conceive of $L$ as a regulator process arising from reflection at a barrier. In particular, we have to weaken the conditions $X_{t}\geq b_{t}$ and $\int_{0}^{t}(X_{s}-b_{s})\,\mathrm{d}L_{s}=0$, characterising the Skorokhod problem underlying the strong formulation of McKean–Vlasov SDE (1.3). This is addressed by the following definition, for which we fix a filtered probability space $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ carrying an $\mathcal{F}_{0}$-measurable random variable $\xi$, an $\mathbb{F}$-Brownian motion $W$, as well as a subfiltration $\mathbb{G}$ of $\mathbb{F}$ independent of $(\xi,W)$. We also denote by $\mathbb{F}^{\xi,W}=(\mathcal{F}^{\xi,W}_{t})_{t\geq 0}$ the filtration given by $\mathcal{F}^{\xi,W}_{t}=\sigma(\xi,W_{s}\mathpunct{:}s\in[0,t])$ for $t\geq 0$ and say that an $\mathcal{M}_{1}([0,\infty)\times\mathbb{R})$-valued random variable is $\mathbb{G}$-adapted if for all $t\geq 0$, the random variable $\beta([0,s]\times A)$ is $\mathcal{G}_{t}$-measurable for $s\in[0,t]$ and $A\in\mathcal{B}(\mathbb{R})$.
@@
As hinted at earlier, it can be shown that $(\beta^{N})_{N\geq 1}$ is tight on $\mathcal{M}_{1}([0,\infty)\times\mathbb{R})$. Moreover, we will prove that for any $i\in\{1,\dots,N\}$, the sequence $(L^{N,i})_{N\geq i}$ is tight on $C([0,\infty))$. Note that we add the subscript $N$ to the process $L^{N,i}$ for emphasis. Then, any limit point of the two sequences is indeed a solution to McKean–Vlasov SDE (1.3) in the sense of Definition 1.3.
@@
Assume that $\mathbb{E}[\xi_{1}]<\infty$. Then, for any $i\in\mathbb{N}$, the sequence $(L^{N,i},\beta^{N})_{N\geq i}$ is tight on the space $C([0,\infty))\times\mathcal{M}_{1}([0,\infty)\times\mathbb{R})$ and any weak limit point is a solution to McKean–Vlasov SDE (1.3) in the sense of Definition 1.3.
@@
McKean–Vlasov SDE (1.3), understood in the sense of Definition 1.3, exhibits pathwise uniqueness. For the unique solution $(L,\beta)$ it holds that $L$ is $\mathbb{F}^{\xi,W}$-adapted and $\beta$ is deterministic. In particular, if $\mathbb{E}[\xi_{1}]<\infty$, then for any $i\in\mathbb{N}$, the sequence $(L^{N,i},\beta^{N})_{N\geq i}$ converges weakly to $(L,\beta)$.
The proof of Theorem 1.5 proceeds via an approximation argument, which exploits that for initial conditions with a sufficiently regular distribution, McKean–Vlasov SDE (1.3) can in fact be solved in the strong sense. A generalised solution to McKean–Vlasov SDE (1.3) for an arbitrary initial condition is then uniquely determined as the limit of an approximating sequence of strong solutions with regular initial conditions. The hardest step in this argument consists in establishing the existence of strong solutions for regular initial conditions. Our construction of such strong solutions relies on an alternative representation of McKean–Vlasov SDE (1.3) through another McKean–Vlasov SDE with reflection. In the course of the derivation of this representation, we shall encounter the so-called supercooled Stefan problem. Before proceeding to this, let us state the existence result for strong solutions of McKean–Vlasov SDE (1.3), starting with suitable assumptions.
@@
We assume that the distribution of $\xi$ has a density given by a càdlàg function with finite total variation that does not vanish at zero.
A variation of Assumption 1.6 naturally arises in the treatment of the related supercooled Stefan problem by Delarue, Nadtochiy & Shkolnikov [11]. We shall see further below how it enters our analysis. Under Assumption 1.6 we have the following existence result.
@@ 1.2 Connection with the Supercooled Stefan Problem
for $x\in(b_{t},\infty)$. Next, let us heuristically derive an appropriate boundary condition at $x=b_{t}$. It follows from the occupation time formula (see e.g. [29, Chapter VI, Corollary 1.6]) that
@@
where $\alpha=\frac{1}{2\gamma}$. This is a moving boundary problem of Stefan-type and appeared in work by Cabezas, Dembo, Sarantsev & Sidoravicius [7] on a scaling limit for the infinite Atlas model. In the infinite Atlas model, we have a semi-infinite collection $(X^{i})_{i\in\mathbb{N}}$ of particles, started from the points of a Poisson process with rate $\lambda>0$. Similar to (1.1), the lowest ranked amongst all particles receives a drift of intensity $\gamma$. Applying the diffusive scaling $X^{N,i}_{t}=\frac{1}{N}X^{i}_{N^{2}t}$, one obtains an infinite version of Equation (1.1). Cabezas, Dembo, Sarantsev & Sidoravicius [7] describe the limit of this system as $N\to\infty$ through a version of PDE (1.7) whose initial condition has constant density $\lambda$ and, consequently, infinite mass. The very recent work by Atar & Budhiraja [2] considers arbitrary locally finite measures as initial densities. We provide a detailed comparison of our results with those of [7] and [2] in Subsection 1.4.
At first glance, PDE (1.7) looks like a classical one-phase Stefan problem with inhomogeneous Dirichlet boundary condition $\mu_{t}(b_{t})=\frac{1}{\alpha}$. As such, it may seem more benign than the supercooled Stefan problem with homogeneous Dirichlet boundary condition, where the dynamics of the boundary $b_{t}$ carry the opposite sign, i.e. $\dot{b}_{t}=\alpha\partial_{x}\mu_{t}(b_{t})$. Indeed, for the supercooled Stefan problem, a homogeneous Dirichlet boundary condition implies that the gradient $\partial_{x}\mu_{t}(b_{t})$ of the solution at the boundary is nonnegative. Hence, the boundary advances towards the particles, which can lead to a critical build-up of mass near the boundary, resulting in a blow-up. Now, if the gradient $\partial_{x}\mu_{t}(b_{t})$ were also nonnegative in PDE (1.7), then, since the boundary dynamics have the opposite sign, the boundary would recede from the particles, so a blow-up would not be expected. However, as $\mu_{t}(b_{t})=\frac{1}{\alpha}>0$ in PDE (1.7), the gradient $\partial_{x}\mu_{t}(b_{t})$ may be both positive and negative, corresponding to the fact that the boundary $b_{t}$ can both increase and decrease. Hence, the positive feedback present in the supercooled Stefan problem may also arise in PDE (1.7). In fact, it turns out that PDE (1.7) is the supercooled Stefan problem in disguise. This point can be made transparent by applying a simple transformation to $\mu_{t}$.
@@
This is precisely the supercooled Stefan problem. Since $\mu_{t}(x)\geq 0$, we have the upper bound $\nu_{t}(x)\leq\frac{1}{\alpha}$. Note, however, that in general, $\nu_{t}(x)$ may be negative. To get a better feeling for PDE (1.8), suppose that $\mu_{0}$ is nonincreasing. Then, $\nu_{0}$ is nondecreasing, $\nu_{0}(x)\geq\nu_{0}(0)=0$ for $x\in(0,\infty)$, and $\lim_{x\to\infty}\nu_{0}(x)=\frac{1}{\alpha}-\lim_{x\to\infty}\mu_{0}(x)=\frac{1}{\alpha}$. The inverse relationship between the (asymptotically attained) bound on the initial condition, namely $\frac{1}{\alpha}$, and the feedback coefficient $\alpha$ puts the equation exactly in the critical regime. In this setting, Baker, Hambly & Jettkant [3] show that PDE (1.8) has a unique global solution in a suitably defined weak sense. In particular, the resulting boundary $b$ is a continuous nondecreasing process. If the feedback were any stronger (while $\lim_{x\to\infty}\nu_{0}(x)$ was kept fixed), then PDE (1.8) would break down at a finite time horizon [3, Theorem 1.12]. The construction of the global solution in the critical regime is based on a probabilistic representation of PDE (1.8) in terms of a reflected McKean–Vlasov SDE. Our goal is to extend this probabilistic representation to the general case, where $\mu_{0}$ is not necessarily nonincreasing. Then, the only information we have on $\nu_{0}$ is that $\nu_{0}(0)=0$, $\nu_{0}(x)\leq\frac{1}{\alpha}$ for $x\in(0,\infty)$, and $\lim_{x\to\infty}\nu_{0}(x)=\frac{1}{\alpha}$. Thus, we are still in the critical regime, but $\nu_{0}$ need not be nonnegative or nondecreasing. Neither (global) existence nor uniqueness is now guaranteed (or even considered) by [3].

### v2

@@ 1 Introduction
The Atlas model (1.1) was introduced by Fernholz in his seminal monograph [undefn] on Stochastic Portfolio Theory (see Example 5.3.3). In this context, the particle states represent the logarithmic market capitalisations of constituents of a large stock market. Since then, there has been a steady interest in finite and infinite versions of the Atlas model [undefc, undefz, undefh, undefaa, undefl, undeff, undefa] and more general systems with coefficients depending on particle ranks. Our goal is to understand the behaviour of the system (1.1) as the number $N$ of particles is taken to infinity. In particular, we wish to derive an SDE that describes the dynamics of a representative particle in the limit $N\to\infty$. Due to the interactive nature of the system, we should expect this SDE to be of McKean–Vlasov type.
@@
for $t\geq 0$. The defining features of the Skorokhod problem in one dimension are recalled in Appendix A, Definition A.1.
@@ 1.1 The Mean-Field Limit
where $\tilde{W}=(\tilde{W}_{t})_{t\geq 0}$ is a Brownian motion and $L^{X^{(2)}-X^{(1)}}=(L^{X^{(2)}-X^{(1)}}_{t})_{t\geq 0}$ denotes the local time at zero of the first gap $X^{(2)}-X^{(1)}$. This formula follows e.g. from [undefn, Proposition 4.1.11], which derives the dynamics of ranked processes. A priori, it is not at all clear that the exploding drift $\gamma Nt$ on the right-hand side of (1.4) should be offset by the local time at zero of the gap $X^{(2)}-X^{(1)}$ in precisely such a way that we wind up with a continuous and deterministic barrier $b$. This motivates us to search for a relaxation of the reflected McKean–Vlasov SDE (1.3), which make sense even when the limit of $(B^{N})_{N\geq 1}$ could in principle be a much rougher object.
@@
Then, as soon as we can get sufficient control on the magnitude of $B^{N}$ as $N\to\infty$, it follows that $(\beta^{N})_{N\geq 1}$ is tight on $\mathcal{M}_{1}([0,\infty)\times\mathbb{R})$. Here $\mathcal{M}_{1}([0,\infty)\times\mathbb{R})$ is the set of measures $m$ on $0,\infty)\times\mathbb{R}$ such that $m([0,t]\times\mathbb{R})=t$ for $t\geq 0$, topologised with the coarsest topology that makes the maps $\mathcal{M}_{1}([0,\infty)\times\mathbb{R})\to\mathcal{P}([0,n]\times\mathbb{R})$, $m\mapsto\frac{1}{n}m|_{[0,n]\times\mathbb{R}}$ for $n\in\mathbb{N}$, continuous. If the sequence $(\beta^{N})_{N\geq 1}$ is tight on $\mathcal{M}_{1}([0,\infty)\times\mathbb{R})$, it admits a subsequential weak limit $\beta$ with values in $\mathcal{M}_{1}([0,\infty)\times\mathbb{R})$. We wish to phrase our generalisation of McKean–Vlasov SDE (1.3) in terms of such $\beta$. As before, the state $(X_{t})_{t\geq 0}$ of the representative particle takes the form $X_{t}=\xi+W_{t}+L_{t}$ for a nondecreasing continuous process $L=(L_{t})_{t\geq 0}$ started from zero. However, we can no longer conceive of $L$ as a regulator process arising from reflection at a barrier. In particular, we have to weaken the conditions $X_{t}\geq b_{t}$ and $\int_{0}^{t}(X_{s}-b_{s})\,\mathrm{d}L_{s}=0$, characterising the Skorokhod problem underlying the strong formulation of McKean–Vlasov SDE (1.3).
This is addressed by the following definition, for which we fix a filtered probability space $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ carrying an $\mathcal{F}_{0}$-measurable random variable $\xi\sim\xi_{1}$, an $\mathbb{F}$-Brownian motion $W$, as well as a subfiltration $\mathbb{G}$ of $\mathbb{F}$ independent of $(\xi,W)$. We also denote by $\mathbb{F}^{\xi,W}=(\mathcal{F}^{\xi,W}_{t})_{t\geq 0}$ the filtration given by $\mathcal{F}^{\xi,W}_{t}=\sigma(\xi,W_{s}\mathpunct{:}s\in[0,t])$ for $t\geq 0$ and say that an $\mathcal{M}_{1}([0,\infty)\times\mathbb{R})$-valued random variable is $\mathbb{G}$-adapted if for all $t\geq 0$, the random variable $\beta([0,s]\times A)$ is $\mathcal{G}_{t}$-measurable for $s\in[0,t]$ and $A\in\mathcal{B}(\mathbb{R})$.
@@
As hinted at earlier, it can be shown that $(\beta^{N})_{N\geq 1}$ is tight on $\mathcal{M}_{1}([0,\infty)\times\mathbb{R})$. Moreover, we will prove that for any $i\in\{1,\dots,N\}$, the sequence $(L^{N,i})_{N\geq i}$ is tight on $C([0,\infty))$. Note that we add the subscript $N$ to the process $L^{N,i}$ for emphasis. Any limit point of the two sequences is a solution to McKean–Vlasov SDE (1.3) in the sense of Definition 1.3.
@@
For any $i\in\mathbb{N}$, the sequence $(L^{N,i},\beta^{N})_{N\geq i}$ is tight on the space $C([0,\infty))\times\mathcal{M}_{1}([0,\infty)\times\mathbb{R})$ and any weak limit point is a solution to McKean–Vlasov SDE (1.3) in the sense of Definition 1.3.
@@
McKean–Vlasov SDE (1.3), understood in the sense of Definition 1.3, exhibits pathwise uniqueness. For the unique solution $(L,\beta)$ it holds that $L$ is $\mathbb{F}^{\xi,W}$-adapted and $\beta$ is deterministic. In particular, for any $i\in\mathbb{N}$, the sequence $(L^{N,i},\beta^{N})_{N\geq i}$ converges weakly to $(L,\beta)$.
The proof of Theorem 1.5 proceeds via an approximation argument, which exploits that for initial conditions with a sufficiently regular distribution, McKean–Vlasov SDE (1.3) can in fact be solved in the strong sense. A generalised solution to McKean–Vlasov SDE (1.3) for an arbitrary initial condition is then uniquely determined as the limit of an approximating sequence of strong solutions with regular initial conditions.
The most challenging step in this argument consists in establishing the existence of strong solutions for regular initial conditions. Our construction of such strong solutions relies on an alternative representation of McKean–Vlasov SDE (1.3) through another McKean–Vlasov SDE with reflection. In the course of the derivation of this representation, we shall encounter the so-called supercooled Stefan problem. Before proceeding to this, let us state the existence result for strong solutions of McKean–Vlasov SDE (1.3), starting with suitable assumptions.
@@
We assume that the distribution of $\xi$ has a density given by a càdlàg function $0,\infty)\to[0,\infty)$ with finite total variation that does not vanish at zero.
A variation of Assumption 1.6 naturally arises in the treatment of the related supercooled Stefan problem by Delarue, Nadtochiy & Shkolnikov [undefj]. We shall see further below how it enters our analysis. Under Assumption 1.6 we have the following existence result.
@@
Remark 1.8.
If $\xi$ follows an exponential distribution with rate $2\gamma$, then the solution $b$ to McKean–Vlasov SDE (1.3) is given by $b_{t}=\gamma t$ for $t\geq 0$. In that case, the distribution $\mathcal{L}(\tilde{X}_{t})$ of the solution $(\tilde{X},L)$ to the Skorokhod problem for $\xi+W-b$ is stationary. Moreover, one easily verifies that $\textup{Exp}(2\gamma)$ is the only initial distribution for which this is true. Indeed, stationarity implies that $\mathbb{E}[\xi]=\mathbb{E}[\tilde{X}_{t}]=\mathbb{E}[\xi]-b_{t}+\gamma t$ for $t\geq 0$, so that $b_{t}=\gamma t$. However, the unique stationary distribution of a reflected Brownian motion on the positive half-line with constant drift $\gamma$ is $\textup{Exp}(2\gamma)$.
We shall not discuss convergence to stationarity here, though we note that one can likely borrow ideas from the proof of [undefb, Theorem 1.16] to establish this.
@@ 1.2 Connection with the Supercooled Stefan Problem
for $x\in(b_{t},\infty)$. Next, let us heuristically derive an appropriate boundary condition at $x=b_{t}$. It follows from the occupation time formula (see e.g. [undefab, Chapter VI, Corollary 1.6]) that
@@
where $\alpha=\frac{1}{2\gamma}$. This is a moving boundary problem of Stefan-type and appeared in work by Cabezas, Dembo, Sarantsev & Sidoravicius [undeff] on a scaling limit for the infinite Atlas model. The infinite Atlas model consists of a semi-infinite collection $(X^{i})_{i\in\mathbb{N}}$ of particles, started from the points of a Poisson process with rate $\lambda>0$. Similar to (1.1), the lowest ranked amongst all particles receives a drift of intensity $\gamma$. Applying the diffusive scaling $X^{N,i}_{t}=\frac{1}{N}X^{i}_{N^{2}t}$, one obtains an infinite version of Equation (1.1). Cabezas, Dembo, Sarantsev & Sidoravicius [undeff] describe the limit of this system as $N\to\infty$ through a version of PDE (1.7) whose initial condition has constant density $\lambda$ and, consequently, infinite mass. The very recent work by Atar & Budhiraja [undefa] considers arbitrary locally finite measures as initial densities. We provide a detailed comparison of our results with those of [undeff] and [undefa] in Subsection 1.4.
At first glance, PDE (1.7) looks like a classical one-phase Stefan problem with inhomogeneous Dirichlet boundary condition $\mu_{t}(b_{t})=\frac{1}{\alpha}$. As such, it may seem more benign than the supercooled Stefan problem with homogeneous Dirichlet boundary condition, where the dynamics of the boundary $b_{t}$ carry the opposite sign, i.e. $\dot{b}_{t}=\alpha\partial_{x}\mu_{t}(b_{t})$. Indeed, for the supercooled Stefan problem, a homogeneous Dirichlet boundary condition implies that the gradient $\partial_{x}\mu_{t}(b_{t})$ of the solution at the boundary is nonnegative. Hence, the boundary advances towards the particles, which can lead to a critical build-up of mass near the boundary, resulting in a blow-up.
Now, if the gradient $\partial_{x}\mu_{t}(b_{t})$ were also nonnegative in PDE (1.7), then, since the boundary dynamics have the opposite sign, the boundary would recede from the particles, so a blow-up would not be expected. However, as $\mu_{t}(b_{t})=\frac{1}{\alpha}>0$ in PDE (1.7), the gradient $\partial_{x}\mu_{t}(b_{t})$ may be both positive and negative, corresponding to the fact that the boundary $b_{t}$ can both increase and decrease. Hence, the positive feedback present in the supercooled Stefan problem may also arise in PDE (1.7). In fact, it turns out that PDE (1.7) is the supercooled Stefan problem in disguise. This point can be made transparent by applying a simple transformation to $\mu_{t}$.
@@
This is precisely the supercooled Stefan problem. Since $\mu_{t}(x)\geq 0$, we have the upper bound $\nu_{t}(x)\leq\frac{1}{\alpha}$. Note, however, that in general, $\nu_{t}(x)$ may be negative. To get a better feeling for PDE (1.8), suppose that $\mu_{0}$ is nonincreasing and bounded from above by $\frac{1}{\alpha}$. Then, $\nu_{0}$ is nondecreasing, $\nu_{0}(0)=\frac{1}{\alpha}-\mu_{0}(0)<\frac{1}{\alpha}$, $0\leq\nu_{0}(x)\leq\frac{1}{\alpha}$ for $x\in[0,\infty)$, and $\lim_{x\to\infty}\nu_{0}(x)=\frac{1}{\alpha}-\lim_{x\to\infty}\mu_{0}(x)=\frac{1}{\alpha}$. The inverse relationship between the (asymptotically attained) bound on the initial condition, namely $\frac{1}{\alpha}$, and the feedback coefficient $\alpha$ driving the boundary puts the equation exactly in the critical regime. In this setting, Baker, Hambly & Jettkant [undefb] show that PDE (1.8) has a unique global solution (in a suitably defined weak sense). In particular, the resulting boundary $b$ is a continuous nondecreasing function. If the feedback were any stronger (while $\lim_{x\to\infty}\nu_{0}(x)$ was kept fixed), then PDE (1.8) would break down at a finite time horizon [undefb, Theorem 1.12].
The construction of the global solution in the critical regime in [undefb] is based on a probabilistic representation of PDE (1.8) in terms of a reflected McKean–Vlasov SDE. Our goal is to extend this probabilistic representation to the general case, where $\mu_{0}$ is not necessarily nonincreasing. Then, the only information we have on $\nu_{0}$ is that $\nu_{0}(0)<\frac{1}{\alpha}$, $\nu_{0}(x)\leq\frac{1}{\alpha}$ for $x\in(0,\infty)$, and $\lim_{x\to\infty}\nu_{0}(x)=\frac{1}{\alpha}$. Thus, we are still in the critical regime, but $\nu_{0}$ need not be nonnegative or nondecreasing. Neither (global) existence nor uniqueness is now guaranteed (or even considered) by [undefb].